# Inexact Bregman Proximal Gradient Method and its Inertial Variant with Absolute and Partial Relative Stopping Criteria

### v4

@@ 3.1 Examples of verifying the stopping criteria
As discussed in [12, Section 3.1], when $\phi$ is of Legendre type (see [61, Page 258]) and $\nabla\phi(\bm{x}^{k})-\lambda^{-1}\left(\nabla f(\bm{x}^{k})+B^{\top}\bm{z}\right)\in\mathrm{dom}\nabla\phi^{*}$, the above infimum is attainable at $p^{k}_{\lambda}\left(\nabla f(\bm{x}^{k})+B^{\top}\bm{z}\right)$, where
@@
We would like to highlight the underlying feasibility difficulty when solving the subproblem in the BPGM, using Example 2. Indeed, when employing the exact BPGM in [12, 50, 69, 74] or the inexact BPGM in [66], one needs to find the exact solution or an inexact solution $\bar{\bm{x}}^{k+1}$ that should be strictly contained in $\mathrm{dom}\,P\cap\mathrm{dom}\,\nabla\phi=\{\bm{x}\in\mathbb{R}^{n}:A\bm{x}=\bm{b},\,\bm{x}>0\}=:\mathrm{relin}\,\mathcal{C}$ (namely, the relative interior of $\mathcal{C}$). However, for most choices of $(A,\bm{b})$, a subsolver for solving (3.5) or its dual (3.6) may only return a candidate approximate solution $\bm{x}^{k,t}$ that is located outside $\mathcal{C}$. Hence, one has to further perform a proper projection/rounding step at $\bm{x}^{k,t}$ to compute an approximate solution $\bar{\bm{x}}^{k+1}$ in $\mathrm{relin}\,\mathcal{C}$, and then check the inexact condition at $\bar{\bm{x}}^{k,t}$. Due to the non-closedness of $\mathrm{relin}\,\mathcal{C}$, this step is indeed difficult to implement especially when some entries of the exact solution are very close to zero. In contrast, as shown in Example 2, our two-point-type inexact condition requires separately $\bm{x}^{k,t}\in\mathcal{X}\cap\mathrm{int}\,\mathrm{dom}\,\phi$ and $\widetilde{\bm{x}}^{k,t}\in\mathrm{dom}\,P\cap\mathrm{dom}\,\phi=\mathcal{C}$, where $\mathcal{X}\supseteq\mathrm{dom}\,P\cap\mathrm{dom}\,\phi$ can be any closed convex set satisfying Assumption A3. The point $\bm{x}^{k,t}$ (which is generally located outside $\mathcal{C}$) can be easily obtained from the subsolver and the point $\widetilde{\bm{x}}^{k,t}$ can be computed by performing a proper projection/rounding step at $\bm{x}^{k,t}$ over $\mathcal{C}$, which is clearly easier and more practical than computing a point in $\mathrm{relin}\,\mathcal{C}$. In this regard, our iBPGM is more favorable than the exact BPGM in [12, 50, 69, 74] and the inexact BPGM in [66]. More discussions on the potential advantages of such a two-point-type inexact condition over existing ones can be found in [73, Section 4].
@@ 3.2 Convergence analysis for the iBPGM
We next establish the convergence of our iBPGM under the (AbSC) or (ReSC) stopping criterion when solving the subproblems. The analysis is inspired by [12, 50], but is more involved due to the approximate minimization of the subproblem under (AbSC) or (ReSC). We first give the following sufficient-descent-like property.
@@
Suppose that Assumption A holds, $\lambda>L$, and $0<\sigma<\frac{\lambda-L}{\lambda}$. Let $\{\bm{x}^{k}\}$ and $\{\widetilde{\bm{x}}^{k}\}$ be the sequences generated by the iBPGM with the stopping condition (ReSC) in Algorithm 1, and let $e(\bm{x}):=F(\bm{x})-F^{*}$, where $F^{*}:=\inf\left\{F(\bm{x}):\bm{x}\in\mathcal{Q}\right\}>-\infty$. Then, the following statements hold.
@@
All or part of the above assumptions are typically utilized in the study of the convergence of the Bregman proximal point algorithm or the Bregman proximal gradient method in the convex setting; see, for example, [12, 27, 28, 35, 36, 65, 67, 73]. These assumptions can be satisfied by several well-known kernel functions, such as the quadratic kernel function $\phi(\bm{x}):=\frac{1}{2}\|\bm{x}\|^{2}$, the entropy kernel function $\phi(\bm{x}):=\sum_{i=1}^{n}x_{i}(\log x_{i}-1)$, and the Hellinger kernel function $\phi(\bm{x}):=-\sum_{i=1}^{n}\sqrt{1-x_{i}^{2}}$. Indeed, Bauschke and Borwein have demonstrated in [13, Proposition 3.3 and Remark 3.4] that Assumption B3 can be satisfied by all separable essentially smooth functions. Moreover, Solodov and Svaiter have shown in [65, Theorem 2.4] that Assumption B4 (convergence consistency) holds under the following conditions: (i) $\phi$ is strictly convex and continuous on $\mathrm{dom}\,\phi$; (ii) $\phi$ is continuously differentiable on $\mathrm{int}\,\mathrm{dom}\,\phi$; (iii) Assumptions B1, B2 and B3 are satisfied. More in-depth discussions and explanations on functions with these properties can be found in [13, 26] and references therein. With Assumption B, we are now ready to present the convergence results for the iterates generated by Algorithm 1.
@@ 4 An inertial variant of the iBPGM
In this section, we shall develop an inertial variant of our iBPGM based on two types of inexact stopping criteria to obtain a possibly faster convergence speed. Before proceeding, we introduce an additional restricted relative smoothness assumption, which modifies the one proposed by Gutman and Peña in [39, Section 3.3]. This assumption is crucial for developing the convergence rate of our inertial variant and can subsume the related conditions used in [10, 40, 68, 69] for developing inertial methods.
@@
while the termination criterion is not met, do
@@
Our v-iBPGM in Algorithm 2 is inspired by Hanzely, Richtárik and Xiao’s accelerated Bregman proximal gradient method [40], which extends Auslender and Teboulle’s improved interior gradient algorithm [10, Section 5] and Tseng’s extension [68, 69] to the relatively smooth setting. However, note that none of the methods in [10, 40, 68, 69] allow the subproblem to be solved approximately. Moreover, the analysis in [40] is established based on the relative smoothness condition and a crucial triangle scaling property (TSP) for the Bregman distance. Since TSP may not imply the strong convexity of the kernel function, the convergence results in [40] actually cannot recover the related results in [10, 68, 69] when $\nabla f$ is $L$-Lipschitz continuous. In contrast, our analysis shall use the restricted relative smoothness condition (2.1) together with the more general condition (4.1) in Assumption C that could be satisfied by the conditions used in [10, 40, 68, 69]. Thus, our subsequent convergence results can readily subsume the related results in [10, 40, 68, 69] when the subproblem is solved exactly.
@@
$+\lambda\theta_{k}^{\gamma-1}\big{(}\mathcal{D}_{\phi}(\widetilde{\bm{z}}^{k+1},\,\bm{z}^{k+1})+\lambda^{-1}\theta_{k}^{1-\gamma}\delta_{k}\big{)}+|\langle\Delta^{k},\,\widetilde{\bm{z}}^{k+1}-\bm{x}\rangle|.$
@@
$P(\widetilde{\bm{z}}^{k+1})+f(\bm{y}^{k})+\langle\nabla f(\bm{y}^{k}),\,\widetilde{\bm{z}}^{k+1}-\bm{y}^{k}\rangle$ (4.8)
@@ 5.1 Implementation details
We follow [73, Section 6.1] to randomly generate the simulated data. Specifically, we first generate two discrete probability distributions $\big{\{}(a_{i},\,\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big{\}}$ and $\big{\{}(b_{j},\,\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big{\}}$. Here, $\bm{a}:=(a_{1},\cdots\!,a_{m})^{\top}$ and $\bm{b}:=(b_{1},\cdots\!,b_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval (0, 1)01(0,\,1)( 0 , 1 ) and further normalized such that $\sum^{m}_{i}a_{i}=\sum^{n}_{j}b_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
@@ Appendix A Proofs of supporting lemmas
Note from the four points identity (2.3) that
$\langle\,\nabla\phi(\bm{x}^{*})-\nabla\phi(\bm{y}),\,\bm{x}-\widetilde{\bm{x}}^{*}\,\rangle=\mathcal{D}_{\phi}(\bm{x},\,\bm{y})-\mathcal{D}_{\phi}(\bm{x},\,\bm{x}^{*})-\mathcal{D}_{\phi}(\widetilde{\bm{x}}^{*},\,\bm{y})+\mathcal{D}_{\phi}(\widetilde{\bm{x}}^{*},\,\bm{x}^{*})$ (A.2)

### v5

@@ 3.1 Examples of verifying the stopping criteria
As discussed in [12, Section 3.1], when $\phi$ is of Legendre type (see [62, Page 258]) and $\nabla\phi(\bm{x}^{k})-\lambda^{-1}\left(\nabla f(\bm{x}^{k})+B^{\top}\bm{z}\right)\in\mathrm{dom}\nabla\phi^{*}$, the above infimum is attainable at $p^{k}_{\lambda}\left(\nabla f(\bm{x}^{k})+B^{\top}\bm{z}\right)$, where
@@
We would like to highlight the underlying feasibility difficulty when solving the subproblem in the BPGM, using Example 2. Indeed, when employing the exact BPGM in [12, 50, 70, 75] or the inexact BPGM in [67], one needs to find the exact solution or an inexact solution $\bar{\bm{x}}^{k+1}$ that should be strictly contained in $\mathrm{dom}\,P\cap\mathrm{dom}\,\nabla\phi=\{\bm{x}\in\mathbb{R}^{n}:A\bm{x}=\bm{b},\,\bm{x}>0\}=:\mathrm{relin}\,\mathcal{C}$ (namely, the relative interior of $\mathcal{C}$). However, for most choices of $(A,\bm{b})$, a subsolver for solving (3.5) or its dual (3.6) may only return a candidate approximate solution $\bm{x}^{k,t}$ that is located outside $\mathcal{C}$. Hence, one has to further perform a proper projection/rounding step at $\bm{x}^{k,t}$ to compute an approximate solution $\bar{\bm{x}}^{k+1}$ in $\mathrm{relin}\,\mathcal{C}$, and then check the inexact condition at $\bar{\bm{x}}^{k,t}$. Due to the non-closedness of $\mathrm{relin}\,\mathcal{C}$, this step is indeed difficult to implement especially when some entries of the exact solution are very close to zero. In contrast, as shown in Example 2, our two-point-type inexact condition requires separately $\bm{x}^{k,t}\in\mathcal{X}\cap\mathrm{int}\,\mathrm{dom}\,\phi$ and $\widetilde{\bm{x}}^{k,t}\in\mathrm{dom}\,P\cap\mathrm{dom}\,\phi=\mathcal{C}$, where $\mathcal{X}\supseteq\mathrm{dom}\,P\cap\mathrm{dom}\,\phi$ can be any closed convex set satisfying Assumption A3. The point $\bm{x}^{k,t}$ (which is generally located outside $\mathcal{C}$) can be easily obtained from the subsolver and the point $\widetilde{\bm{x}}^{k,t}$ can be computed by performing a proper projection/rounding step at $\bm{x}^{k,t}$ over $\mathcal{C}$, which is clearly easier and more practical than computing a point in $\mathrm{relin}\,\mathcal{C}$. In this regard, our iBPGM is more favorable than the exact BPGM in [12, 50, 70, 75] and the inexact BPGM in [67]. More discussions on the potential advantages of such a two-point-type inexact condition over existing ones can be found in [74, Section 4].
@@ 3.2 Convergence analysis for the iBPGM
We next establish the convergence of our iBPGM under the (AbSC) or (ReSC) when solving the subproblems. The analysis is inspired by [12, 50], but is more involved due to the approximate minimization of the subproblem under (AbSC) or (ReSC). We first give the following sufficient-descent-like property.
@@
Suppose that Assumption A holds, $\lambda>L$, and $0<\sigma<\frac{\lambda-L}{\lambda}$. Let $\{\bm{x}^{k}\}$ and $\{\widetilde{\bm{x}}^{k}\}$ be the sequences generated by the iBPGM with (ReSC) in Algorithm 1, and let $e(\bm{x}):=F(\bm{x})-F^{*}$, where $F^{*}:=\inf\left\{F(\bm{x}):\bm{x}\in\mathcal{Q}\right\}>-\infty$. Then, the following statements hold.
@@
All or part of the above assumptions are typically utilized in the study of the convergence of the Bregman proximal point algorithm or the Bregman proximal gradient method in the convex setting; see, for example, [12, 27, 28, 35, 36, 66, 68, 74]. These assumptions can be satisfied by several well-known kernel functions, such as the quadratic kernel function $\phi(\bm{x}):=\frac{1}{2}\|\bm{x}\|^{2}$, the entropy kernel function $\phi(\bm{x}):=\sum_{i=1}^{n}x_{i}(\log x_{i}-1)$, and the Hellinger kernel function $\phi(\bm{x}):=-\sum_{i=1}^{n}\sqrt{1-x_{i}^{2}}$. Indeed, Bauschke and Borwein have demonstrated in [13, Proposition 3.3 and Remark 3.4] that Assumption B3 can be satisfied by all separable essentially smooth functions. Moreover, Solodov and Svaiter have shown in [66, Theorem 2.4] that Assumption B4 (convergence consistency) holds under the following conditions: (i) $\phi$ is strictly convex and continuous on $\mathrm{dom}\,\phi$; (ii) $\phi$ is continuously differentiable on $\mathrm{int}\,\mathrm{dom}\,\phi$; (iii) Assumptions B1, B2 and B3 are satisfied. More in-depth discussions and explanations on functions with these properties can be found in [13, 26] and references therein. With Assumption B, we are now ready to present the convergence results for the iterates generated by Algorithm 1.
@@ 4 An inertial variant of the iBPGM
In this section, we shall develop an inertial variant of our iBPGM based on two types of inexact stopping criteria to obtain a possibly faster convergence rate. Before proceeding, we introduce an additional restricted relative smoothness assumption, which modifies the one proposed by Gutman and Peña in [39, Section 3.3]. This assumption is crucial for developing the convergence rate of our inertial variant and can subsume the related conditions used in [10, 40, 69, 70] for developing inertial methods.
@@
while a termination criterion is not met, do
@@
Our v-iBPGM in Algorithm 2 is inspired by Hanzely, Richtárik and Xiao’s accelerated Bregman proximal gradient method [40], which extends Auslender and Teboulle’s improved interior gradient algorithm [10, Section 5] and Tseng’s extension [69, 70] to the relatively smooth setting. However, note that none of the methods in [10, 40, 69, 70] allow the subproblem to be solved approximately. Moreover, the analysis in [40] is established based on the relative smoothness condition and a crucial triangle scaling property (TSP) for the Bregman distance. Since TSP may not imply the strong convexity of the kernel function, the convergence results in [40] actually cannot recover the related results in [10, 69, 70] when $\nabla f$ is $L$-Lipschitz continuous. In contrast, our analysis shall use the restricted relative smoothness condition (2.1) together with a more general condition (4.1) in Assumption C that could be satisfied by the conditions used in [10, 40, 69, 70]. Thus, our subsequent convergence results can readily subsume the related results in [10, 40, 69, 70] when the subproblem is solved exactly.
@@
$~{}~{}+\lambda\theta_{k}^{\gamma-1}\big{(}\mathcal{D}_{\phi}(\widetilde{\bm{z}}^{k+1},\,\bm{z}^{k+1})+\lambda^{-1}\theta_{k}^{1-\gamma}\delta_{k}\big{)}+|\langle\Delta^{k},\,\widetilde{\bm{z}}^{k+1}-\bm{x}\rangle|.$
@@
$\quad P(\widetilde{\bm{z}}^{k+1})+f(\bm{y}^{k})+\langle\nabla f(\bm{y}^{k}),\,\widetilde{\bm{z}}^{k+1}-\bm{y}^{k}\rangle$ (4.8)
@@ 5.1 Implementation details
We follow [74, Section 6.1] to randomly generate the simulated data. Specifically, we first generate two discrete probability distributions $\big{\{}(a_{i},\,\bm{p}_{i})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:i=1,\cdots,m\big{\}}$ and $\big{\{}(b_{j},\,\bm{q}_{j})\in\mathbb{R}_{+}\times\mathbb{R}^{3}:j=1,\cdots,n\big{\}}$. Here, $\bm{a}:=(a_{1},\cdots\!,a_{m})^{\top}$ and $\bm{b}:=(b_{1},\cdots\!,b_{n})^{\top}$ are probabilities/weights, which are generated from the uniform distribution on the open interval (0, 1)01(0,\,1)( 0 , 1 ) and further normalized such that $\sum^{m}_{i}a_{i}=\sum^{n}_{j}b_{j}=1$. Moreover, $\{\bm{p}_{i}\}$ and $\{\bm{q}_{j}\}$ are support points whose entries are drawn from a five-component multivariate Gaussian mixture distribution, with a mean vector $(-20,10,0,10,20)^{\top}$ and a variance vector $(5,5,5,5,5)^{\top}$, using randomly generated weights. Then, the cost matrix $C$ is generated by $c_{ij}=\|\bm{p}_{i}-\bm{q}_{j}\|^{2}$ for $1\leq i\leq m$ and $1\leq j\leq n$ and normalized by dividing (element-wise) by its maximal entry.
@@ Appendix A Proofs of supporting lemmas
Note from the four-points identity (2.3) that
$\langle\,\nabla\phi(\bm{x}^{*})-\nabla\phi(\bm{y}),\,\bm{x}-\widetilde{\bm{x}}^{*}\,\rangle=\mathcal{D}_{\phi}(\bm{x},\,\bm{y})-\mathcal{D}_{\phi}(\bm{x},\,\bm{x}^{*})-\mathcal{D}_{\phi}(\widetilde{\bm{x}}^{*},\,\bm{y})+\mathcal{D}_{\phi}(\widetilde{\bm{x}}^{*},\,\bm{x}^{*}).$ (A.2)